OpenStudy (lilday12):

Evaluate the expression below. -2-(-6)

OpenStudy (anonymous):

- and - make a + so now just do -2+6 which is 6-2=?

OpenStudy (lilday12):

It 4

OpenStudy (jjcoolja12):

−2−(−6) =−2−(−6) =−2+6 =4